### v2

# Bringing Diversity from Diffusion Models to Semantic-Guided Face Asset Generation

Abstract.
High-quality 3D face asset creation remains costly due to reliance on controlled capture setups and manual processing, limiting scalability and diversity. We introduce a fully automated, semantically controllable framework for generating PBR-ready 3D facial assets without requiring dedicated scans. Our pipeline begins with a diffusion-based data synthesis stage, where 2D portrait samples from a pre-trained diffusion model are converted into 44K textured 3D face reconstructions via our proposed geometry recovery and texture normalization algorithm, which aligns arbitrarily shaded outputs into clean albedo space. Using this dataset, we train a disentangled adversarial generator that maps semantic attributes (age, gender, ethnicity) to UV-space geometry and albedo, enabling both direct sampling and continuous latent editing while preserving identity. A refinement stage further produces PBR materials and secondary assets (eyeballs, teeth, gums). The resulting system supports controllable face generation and post-editing in real time and exports directly to standard rendering and animation pipelines. We evaluate each component extensively and provide a web-based interactive interface to showcase practical deployment.
1. Introduction
Creating realistic 3D human faces has long been a central goal in both industry and academia. A high-quality 3D face asset integrates detailed skin textures, accurate geometry, and material properties to function within a physically based rendering (PBR) pipeline. Demand for such assets continues to surge across gaming, film, teleconferencing, and AR/VR applications. However, producing them remains labor-intensive and expensive—typically requiring skilled artists and, in the case of scans, access to a carefully calibrated facial capture studio. Even today, generating production-ready 3D face assets is still a costly and time-consuming endeavor.
Despite extensive efforts to automate high-quality 3D face creation for broader accessibility, current face generation models remain constrained by the scarcity and imbalance of available facial asset datasets. Industrial systems like MetaHuman (Fang et al., 2021) rely on limited preset assets, while traditional 3D Morphable Models (Paysan et al., 2009; Booth et al., 2018; Cao et al., 2014; Li et al., 2017) and learning-based approaches (Li et al., 2020b; Yang et al., 2020; Zhang et al., 2023a) struggle to jointly model geometry and texture, failing to capture the full diversity of real human faces or offer precise semantic control. Data limitations are especially evident in facial texture generation—skin tone, freckles, wrinkles, and pores are poorly represented, leading to biased outputs such as DreamFace (Zhang et al., 2023a), which predominantly synthesizes Asian-like textures due to dataset skew (over 90% of FaceScape (Yang et al., 2020) is Asian). Recent methods that leverage CLIP (Radford et al., 2021) or large diffusion models improve diversity through knowledge distillation, as seen in AvatarClip (Hong et al., 2022) and DreamAvatar (Cao et al., 2023), while LucidDreamer (Liang et al., 2023) enhances detail using Interval Score Matching. However, these approaches still suffer from artifacts, incomplete assets, slow inference, and lack of animation-ready outputs, leaving a significant gap between generative quality and production usability.
Our approach addresses data scarcity in 3D face asset generation by leveraging large diffusion models for diverse synthesis while introducing three key innovations that bridge the gap between scanned-quality assets and synthetic outputs. (1) Diffusion-generated UV textures are often incomplete and inconsistently shaded, making them unsuitable for PBR rendering. We introduce a texture completion and normalization pipeline to produce clean, fully albedo-compatible maps. (2) We enable precise semantic control and editing by structuring attributes around demographic labels (ethnicity, gender, age) and adopting a disentangled GAN-based framework, which preserves identity during editing more effectively than diffusion-based inversion. (3) We drastically reduce generation time by distilling knowledge into a GAN, achieving 0.014s versus 40s per face compared to a diffusion model on a single A6000 GPU—making high-quality face asset generation both controllable and real-time.
Our training pipeline consists of two major stages. The first stage leverages an image-based diffusion network and a 3D face reconstruction network to create a dataset of diverse, high-quality 3D face assets. As the direct image output of the diffusion network contains lighting and other external visual effects, albedo maps are computed from a proposed texture normalization algorithm. In addition, we apply a sanity check to ensure quality and discard any data with artifacts or mismatched labels, and convert the conditioning text used to generate the data into three controlling attributes: age, gender, and ethnicity. This process resulted in a dataset of 44K 3D face models. The second stage extends from DisUnknown (Xiang et al., 2021) to train a GAN with the processed training data from the first stage. As described, this stage produces the final model that generates 3D face assets consisting of 4K geometry, albedo, specular and displacement maps in the UV space. Thanks to our unique design on adversarial training, the flexible model can either create 3D assets from an attribute description (“generate a 40-year-old, Hispanic male”), or perform inversion on a given image and subsequently edit and reconstruct a 3D asset from the image, based on an attribute description (“turn a face photo into a 3D face of a 40-year-old, Hispanic male, without losing their identity”).
We summarize our contributions as follows:
We’ve introduced a comprehensive, practical and novel framework for generating high-quality face assets. This system uses user-defined semantics and attributes to create PBR-based face assets, including base geometry, albedo, specular, and displacement maps, as well as secondary assets such as eyeballs, teeth, and gums. The system also allows for post-generation editing in both geometry and texture, while preserving identity. The generated face avatars can be seamlessly integrated into downstream applications for rendering and animation. Additionally, we’ve developed an interactive web UI for users to explore these features.
We have developed a large, high-quality 3D face database containing albedo/geometry pairs, complete with age, gender, and ethnicity labels. This database exemplifies an effective way to use a pre-trained diffusion model in an industry production pipeline.
We also tackled the challenging problem of domain transfer with unbalanced data amounts in two domains. Our texture normalization framework transfers 44,000 unconstrained images into a domain containing 200 images. This allows us to combine the diversity of one domain with the quality from another. This could inspire research in this field.
2. Related Work
2.1. 3D Face Morphable Model
The 3D Morphable Model (3DMM), as the core component of conventional 3D face generation, was first introduced in (Blanz and Vetter, 1999) as a compact representation of face models in parametric space. Since then, it has been extensively applied in face recognition (Paysan et al., 2009), face reconstruction (Bas et al., 2016; Thies et al., 2016; Gecer et al., 2019), and avatar creation (Nagano et al., 2018; Li et al., 2020c; Ghafourzadeh et al., 2020). A comprehensive overview of 3DMM is provided in (Egger et al., 2020). Efforts have been made to improve the expressiveness, quality, and parameterization capability of 3DMM over the past 20 years.
Previous work (Paysan et al., 2009; Cao et al., 2014; Booth et al., 2016; Li et al., 2017) reduced the cost and labor required for data acquisition and registration. (Paysan et al., 2009) introduced the first publicly available 3DMM, while (Booth et al., 2018) developed a more diverse linear model using approximately 10,000 scans. For expression modeling, blendshapes were introduced to a bilinear model (Cao et al., 2014; Li et al., 2017). Deep learning methods have since enhanced 3DMM capabilities, enabling more diverse and robust representations (Abrevaya et al., 2018; Ranjan et al., 2018; Dai et al., 2020; Li et al., 2020a; Smith et al., 2020; Chandran et al., 2020). (Yang et al., 2020; Li et al., 2020d) constructed high-quality 3DMM with pore-level resolution data, and (Yang et al., 2020) provided a large-scale textured 3D face dataset that represents geometry through both rough shapes and detailed displacement maps. (Li et al., 2020d) developed a non-linear 3DMM using high-resolution face scans, incorporating material attributes for physically-based rendering. However, challenges remain in expensive data collection and limited diversity.
When applied to 3D face generation, conventional 3DMMs such as (Paysan et al., 2009; Booth et al., 2018; Cao et al., 2014; Li et al., 2017) offer parameterized models for face modeling. However, they lack explicit control over individual attributes and struggle to effectively integrate textures with underlying geometry. This limitation restricts their usefulness in applications requiring customized human face generation. While learning-based 3DMMs like (Li et al., 2020b; Yang et al., 2020) enable joint modeling of texture and geometry, these models still cannot provide adequate semantic or attribute control during generation.
2.2. Text-to-3D Face
Advancements in text-to-3D avatar generation (Michel et al., 2022; Hong et al., 2022; Wu et al., 2023) have leveraged the vision-language model CLIP (Radford et al., 2021) to map natural language descriptions into latent spaces, enabling the generation of 2D images that are then converted into 3D avatars. While this approach allows attribute and semantic control, it often suffers from texture artifacts and inconsistent multi-view outputs. Beyond avatar generation, text-to-3D content creation in general has rapidly evolved, with recent works (Cao et al., 2023; Chen et al., 2023a; Lin et al., 2023; Metzer et al., 2023; Poole et al., 2022) excelling at generating diverse 3D models from text. These methods employ pre-trained text-to-image diffusion models as priors to guide the training of parameterized 3D models, ensuring multi-view consistency and text alignment through Score Distillation Sampling (SDS). However, SDS encourages average-seeking behaviors in the 3D representation, leading to over-smoothed results. LucidDreamer (Liang et al., 2023) addresses this issue with Interval Score Matching (ISM), improving detail quality, but their method remains computationally expensive — taking approximately 36 minutes to generate a 3D model. A challenge in text-to-3D generation is that the appearance often entangles reflectance and illumination. Intrinsic decomposition addresses this issue by separating reflectance (albedo) from lighting, and diffusion priors (Chen et al., 2024; Luo et al., 2024; Kocsis et al., 2024; Li et al., 2025) have recently been used to improve such decomposition in a general image. At the same time, text-to-3D face avatar generation requires producing a usable 3D facial asset, which imposes additional constraints than those for general images.
Among existing methods, DreamFace (Zhang et al., 2023a) has made significant progress in generating high-quality face avatars. It separately trains a geometry and appearance model, fine-tuning a generic diffusion model with 618 textures from sources like FaceScape (Yang et al., 2020) and 3DScanStore (3DScanStore, 2023). However, several limitations exist:1.) Limited geometric diversity – Geometry is initialized from a 3DMM with only 100 bases. 2.) Insufficient and biased texture data – 618 textures are inadequate for fine-tuning a diffusion model trained on vast in-the-wild datasets, and FaceScape’s 90% Asian demographic introduces bias. 3.) Decoupled geometry and texture generation – Ignoring their correlation, as highlighted in (Li et al., 2020b). 4.) Challenging inversion in diffusion models – Editing (e.g., tattoos, makeup) is constrained to one-pass modifications without precise control. 5.) Lack of quality control – Outputs are inconsistent, and quality depends heavily on user prompts.
Inspired by DreamFusion (Poole et al., 2022), LucidDreamer (Liang et al., 2023) and HeadStudio (Zhou et al., 2024b), we aim to refine generative 3D generation. While these models generate 3D faces from text, they often produce artifacts and require manual quality control, making them unsuitable for direct application. Nonetheless, pre-trained image diffusion models, trained on large datasets, naturally enhance diversity, provide annotations and offer great accessibility. Our approach harnesses the rich priors of image diffusion models to improve quality, control, and usability in 3D avatar generation.
2.3. Semantic Face Generation and Manipulation
Generative Adversarial Networks (GANs), pioneered by (Goodfellow et al., 2014), have been well studied in the setting of semantic editing. A key challenge lies in achieving semantic and disentangled control in generative models, e.g., randomly changing one specific attribute while preserving the other attributes. Efforts to interpolate in the latent space for smooth output variation have been explored by (Laine, 2018; Shao et al., 2018), while recent research focuses on disentangling the latent space for semantic control (Shen et al., 2020; Härkönen et al., 2020; Shen and Zhou, 2020; Jiang et al., 2021; Zheng et al., 2021). Observations of vector arithmetic in latent space by (Radford et al., 2016; Upchurch et al., 2017) have led to unsupervised disentanglement methods. Semi-supervised methods apply principal component analysis for attribute identification (Härkönen et al., 2020), while supervised methods use labeled data for latent space factorization (Shen et al., 2020; Kowalski et al., 2020). Alternative approaches utilize 3DMM for semantic control over pre-trained StyleGAN latent space (Tewari et al., 2020a; Wang et al., 2022) or train unsupervised data with labels from attribute classifiers (Khodadadeh et al., 2022). Also, to transfer such controllability to real image editing, the GAN inversion methods (Abdal et al., 2019; Zhu et al., 2020; Tewari et al., 2020b) propose to map the real image into the latent space and thus can manipulate real images. Recently, (Fadaeinejad et al., 2025) incorporated artist input into a GAN-based framework, enabling greater controllability in the avatar creation workflow.
2.4. Facial Texture Synthesis
3D facial texture synthesis has evolved from traditional geometric and photometric methods to advanced neural network-based techniques. Traditional photometric methods, relying on polarized light to capture skin reflectance properties, laid the groundwork for detailed texture mapping (Debevec et al., 2000; Ghosh et al., 2011). Innovations then made high-quality, single-shot captures possible for both facial geometry and reflectance (Riviere et al., 2020; Lattas et al., 2022).
The advent of deep learning has revolutionized texture synthesis. Neural networks are employed for photorealistic textures, combining low and high-frequency methods (Saito et al., 2017; Yamaguchi et al., 2018; Chen et al., 2019; Huynh et al., 2018). Generative adversarial networks (GANs) have further enhanced the generation of detailed and realistic textures, using patch-based approaches and focusing on physical texture properties like diffuse and specular albedos (Gecer et al., 2019; Lattas et al., 2020, 2021; Dib et al., 2021; Lattas et al., 2023). Recent developments employ the denoising diffusion probabilistic models and text inputs for texture creation (Zhang et al., 2023a). Super-resolution techniques enable the generation of high-resolution textures from lower-resolution inputs, enhancing detail of mesh without altering its geometry (Chen et al., 2023b).
These advancements underscore the goal towards both visual realism and adherence to the principles of physical-based rendering, which guarantees accurate material properties under diverse lighting conditions.
3. Method
Our goal is to create a 3D avatar synthesis model that uses facial attributes (ethnicity, gender and age) to produce high-quality 3D face assets, including 3D geometry and PBR-based material properties such as albedo, specular and displacement maps. To achieve our goal, we start by constructing a large, high-quality and diverse 3D face assets dataset associated with corresponding annotations using a pre-trained diffusion model. We introduce a unique framework in Section 3.1 to construct a 3D face asset dataset that is clean and complete with corresponding labels. Then, a dedicated generative network, as detailed in Section 3.2, is trained using the obtained dataset to create the basic geometries and albedo textures. Lastly, we apply post-processing, as described in Section 4.1, to refine the generated face geometries and albedo textures and generate the specular and displacement maps to complete the PBR assets.
3.1. Data Preparation
We leverage the rich priors of an image diffusion model by using it to synthesize a dataset of high-quality 3D faces with annotations, which is used to train our Text2Avatar 3D face asset generator. This section details three main steps needed to create a dataset of attribute-controlled 3D assets (shapes and albedo textures). Section 3.1.1 describes the generation of 2D portraits that leverages information from both the rich priors of a large image diffusion model (Rombach et al., 2022; Zhang et al., 2023b) and a dataset of scanned 3D faces from LightStage (Li et al., 2020b) and the TrippleGangers dataset (Triplegangers, 2021). Section 3.1.2 describes the construction of 3D face geometries and complete UV textures from the generated portraits. Section 3.1.3 details a normalization method that removes lighting and other external visual effects from the textures to create the clean albedo maps.
3.1.1. Portraits Generation using Diffusion Models
We utilize the pre-trained text-to-image Stable Diffusion v1.5 (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) (trained on normal maps) to generate a rich collection of 2D portraits. Specifically, the diffusion model is given various text descriptions of facial attributes. To avoid ambiguity caused by the language model and ensure precise control, we define three major demographic attributes that contribute most to human appearances and shapes: ethnicity, gender, and age. We also include generic prompts for illumination and quality control, like resolution and framing. For example, ”East Asian female age 20” is one set of demographic attribute combinations. Age groups span from 15 to 75 years old. For gender, we use the labels male, female, and unisex (a gender-neutral semantic category positioned between male and female). We also consider ethnicity/race and geographic location, resulting in 14 distinct categories: African, East Asian, Southeast Asian, South Asian, Middle Eastern, Caucasian, Germanic, Celtic, Slavic, Romance, Australian, Native Americans, Aboriginal, and Pacific Islander. In addition, we condition the portrait generation with normal maps, rendered from a randomly sampled face geometry at the frontal view, from the scan dataset . The normal maps provide head pose and geometry guidance to constrain the synthesized image.
Since random sampling may create a conflict between the desired attributes and the selected geometry, we further reconstruct a refined face geometry based on the alignment between the generated portrait and the sampled face geometry using ReFA (Liu et al., 2022) (see Section 4.1 for details).
In total, 65K portraits with 3D geometry are created at this stage. Ethnicity is uniformly sampled from 14 categories. As for gender, the distribution consists of male, female, and unisex. Age is uniformly sampled from the following groups: , , , , , , , , , , , , and .
3.1.2. Texture Completion
The first step produces a collection of 2D portraits, each with the corresponding attribute labels and predicted geometry. To process this data into 3D assets, we complete the UV-space texture maps from these outputs. Specifically, the texture map is initialized by directly projecting the 2D portraits based on the predicted geometry. However, since only the frontal part of the face is visible, we propose a completion method to fill in the missing regions (e.g. ears) in the boundary of the UV maps. We first collect a dataset of skin textures by baking randomized lighting into the UV-space textures of the scan dataset (see Figure 2). Then, given the projected partial textures, we search for the nearest neighbor with the highest similarity to the projected portrait in , using Peak Signal-to-Noise Ratio (PSNR) as our measure. The complete UV-space texture is then obtained by blending the projected partial textures and the retrieved texture maps with the pyramid blending algorithm (Burt and Adelson, 1983).
Finally, to ensure the quality of the completed textures, a sanity check is applied to filter out results with artifacts to maintain high data quality. The filtering step keeps approximately UV textures from the original portraits. However, these textures still contain baked-in lighting, which is not part of our desired clean albedo. Therefore, a further normalization step is proposed in the next section.
3.1.3. Texture Data Normalization
The goal of this step is to remove lighting and other external visual effects, such as makeup, facial hair, glasses, and shadows, from the UV textures created in the last step. The normalized texture can be considered a clean albedo map that is a part of the output 3D face asset. FFHQ-UV (Bai et al., 2023) is the latest work that can normalize face texture into an albedo image, but we observe that it still struggles to fully preserve identities and facial attributes when applied to our synthesized portraits (see Table 2). We therefore propose a more effective normalization method, by posing texture normalization in this context as a domain transfer problem from the source domain, which consists of synthetic textures under different lighting as shown in Figure 3 (b), to the target domain, which consists of clean albedo as shown in Figure 3 (a). Since the target domain contains only 200 identities while the source domain contains around 44K identities, a convolutional image-to-image translation model directly trained on the whole image would face a high risk of overfitting. The following text describes several designs to mitigate this issue.
First, we assume that lighting affects the appearance of the face approximately uniformly in a reasonably-sized patch. Therefore we divide an input image into a grid of patches and assign a spatially varying factor to each grid corner to account for factors that are independent of albedo. is bilinearly interpolated in the interior of each patch. Under this formulation, image translation is formulated as a function parameterized by an MLP network that takes the image as input and translates each pixel independently based on the spatially-varying factor . However, the image translation is preceded by a patch parameter estimation network, a convolutional network that takes the image as input and computes at patch corners. Effects such as subsurface scattering and inaccuracies due to the non-physical nature of the input images can also be handled by the spatial variation of and do not need to be considered separately.
In addition, different combinations of albedo and lighting can result in the same observed color, so without knowing the true lighting, the albedo is ambiguous. To remove this ambiguity, we compute a skin color by taking the average pixel value over manually selected flat regions of the input face (cheeks and forehead), which is equivalent to a masked average of the image . This value is then provided to the patch parameter estimation network explicitly. Together, the patch parameter estimation network and the MLP translation network are termed the normalization model , which is visualized in Figure 5.
The loss function for training the texture normalization model is constructed as where is the set of scanned albedo textures. During training, we additionally augment the input with albedo images that do not contain lighting and optimize the network to produce identical output to the input in this situation.
The overall training procedure is shown in Figure 6. As commonly done in unpaired image-to-image translation, a patch-based discriminator is employed to ensure that the translated image is in the target domain, trained with the binary cross entropy loss. For adversarial training, the distribution of skin color of the normalized synthesized textures is optimized to match that of the scanned textures, so when normalizing a synthesized texture, the skin color is drawn randomly from the set of skin colors of the scanned textures. Let be the discriminator and the set of synthetic textures. The Discriminator’s loss and the normalization model’s adversarial loss are:
Additionally, we need to ensure that when normalizing a synthesized texture, the output does have the same skin color as the provided skin color:
and are then trained using:
3.2. Base Geometry and Albedo Generation
As mentioned in the introduction, we aim to learn a geometry and albedo texture generation model that can 1) effectively disentangle labeled attributes and unlabeled information (e.g. identity), 2) provide an easy way to perform image inversion, and 3) generate efficiently. With the prepared dataset (geometry, albedo texture with labeled attributes), we propose a two-step generative model, extending from the two-step approach in (Xiang et al., 2021), that satisfies all previous requirements, and we show in Section 4.2.3 that the design is more effective compared to the common practices of generation directly conditioning on the given attributes. The overall procedure is shown in Figure 7.
3.2.1. Learning unlabeled information.
In the first step, a GAN with an autoencoder , a generator and a discriminator is designed to disentangle unlabeled information from the labels (in our case, the gender, age and ethnicity attributes). is first used to extract unlabeled information from the images and discards labeled information. The code computed by the encoder therefore describes information that is independent of the face attributes (e.g. identity information).
How can be trained to disentangle information from attribute labels? If it does discard the labeled information completely, then the conditional distribution of the code, on any value of the attributes should be the same. Based on this observation, we can achieve label disentanglement by using an adversarial classifier that classifies the encoder’s output by their label, and the encoder is trained to make the classifier fail. However, while the conditional distribution of the code is learned to be the same given any label, this distribution has no closed form and cannot be easily sampled. We would like to generate new images, so this conditional distribution must be identical to some known, simple prior. So, different from (Xiang et al., 2021), the code classifier is replaced with a conditional code discriminator.
Given attribute values as conditions, the code discriminator learns to distinguish between the prior distribution and the codes computed by the encoder. We choose the normal distribution as the prior. Let be the attribute label of a training image , and be the training dataset, which are pairs of albedo and geometry:
will also need to retain all unlabeled information. This is achieved using a reconstruction loss. It is worth noting that the first-step generator will not be used as the generator in our finished texture generation model, as it only assists in training the encoder. In addition, takes the labels as an input, since labeled information is removed from the codes . The overall reconstruction loss is:
, and are then trained using:
3.2.2. Training the label-conditioned generator.
As mentioned previously, the trained from step 1 does not model the marginal distribution of the attribute labels as it requires unlabeled information. In the second step, a separate conditional generator is trained to generate base geometry and albedo maps from attributes and a sampled random code . In this step, the trained autoencoder is frozen. A new discriminator receives the unlabeled code as a condition and discriminates whether the generated image has preserved the unlabeled information provided by the code. Specifically, a “real” sample is a triplet consisting of a training image , its code and its label , while a “fake” sample is a triplet consisting of a random code , random labels which are produced by taking the labels of a random training image, and , the image generated using the code and the labels:
and the training procedure of the second step is a plain GAN:
Through experiments, we noticed that although concatenating the one-hot attribute labels to the fully connected part of the discriminator worked as intended, incorporating the unlabeled code in the same way did not give satisfactory results. We speculate that the difficulty lies in the observation that the unlabeled code is much longer than the attribute labels and has a strong spatial structure which the discriminator had to learn from scratch. Therefore, we help the discriminator by generating an image from the label and code it receives using the first-step generator and concatenating its output with the input image, so that the spatial structure is provided by the condition image and need not be re-learned. The detailed architecture of our discriminator is shown in the inset in Figure 7. remains frozen and is not updated along with other parameters of the discriminator.
4. Experiments
4.1. Implementation Details
Synthetic Training Data Generation.
To generate the portraits Figure 2, we use the pre-trained Stable Diffusion v1.5 (Rombach et al., 2022) as the base LDM model and use the ControlNet checkpoint pre-trained on normal maps (Zhang, 2023) to add conditional controls for our frontal view normal map. The sampling method is DDIM with 40 time steps. Classifier-Free Guidance (Ho and Salimans, 2022) scale is 9.0. The output portrait resolution from the LDM model is 1024 1024.
Single-View Face Reconstruction.
In this paper, we perform single-view 3D face reconstruction to predict the 3D geometry of the generated portraits. To train the model, we rendered around synthetic frontal view portraits under different illuminations by using assets in our 3D high-quality database. We combined these with captured multi-view real data. A variant of ReFA (Liu et al., 2022) was used as the baseline to train the model, where the network is modified to reconstruct from a single-view image. The camera optimization is kept fixed during training, as the ground truth pose is known. Therefore, in each network step, only the position map , which corresponds to the 3D geometry of the face, is updated.
The training of the reconstruction model is performed on NVIDIA A100 graphics cards. The network parameters are randomly initialized and are trained using the Adam optimizer for 120,000 iterations with a learning rate set to . For the recurrent face geometry optimizer, we set the inference step to , the grid resolution to , the search radius to .
An example of the refinement to the initial geometry is shown in Figure 8, the refined geometry (c) more accurately matches the portrait (b) compared to the initial geometry (a), with deeper-set eyes that are correctly positioned, a mouth that is properly sized and shaped, more correctly aligned cheekbones, and a chin that reflects the correct depth and contour, resulting in an overall head shape that aligns more closely with (b).
Texture Normalization.
The Patch Parameter Estimation Network is a convolutional network, consisting of alternating kernel 3, stride 1 convolutions and kernel 4, stride 2 convolutions. The number of output channels starts at 16 and doubles after each stride 2 layer. Once the spatial size reduces to 64, the skin color code is broadcast to each location, and two additional kernel 1, stride 1 layers are added. The final layer, which outputs , has 8 output channels. The Pixel Translation Network is an MLP with 6 layers and 64 features in the hidden layers.
To ensure that the discriminator focuses on only the lighting, its capacity is purposefully limited: it consists of an MLP with 6 layers and 64 features in the hidden layers. It takes patches as input and first downsamples them to so that the number of input features is .
Generator
We adopt the StyleGAN2 (Karras et al., 2020) architecture for our generator as well as the convolution part of our discriminator. We also adopt the gradient penalty, path penalty and augmentation from its training procedure.
Asset Refinement
We apply an asset refinement process that augments resolution and supplements missing components required for physically-based rendering. This process consists of the following three stages: First, a super-resolution network (Chen et al., 2023b) is used to upsample the 1K albedo map to 4K, recovering high-frequency skin details. Second, a texture translation network (Liang et al., 2021) predicts corresponding specular and displacement maps from the 4K albedo. The network is trained on our high-quality scanned dataset. To generate displacement maps, the albedo is converted to Lab color space and only the L channel is used as input, isolating geometric features from skin color. We also applied a filter that smooths out the edge of the displacement maps. Finally, we integrate secondary assets such as eyeballs, teeth, and gums, along with predefined Blendshapes, into the base mesh to support animation and rendering.
Performance.
Our base geometry and albedo generation network generates results in 0.014 seconds per subject. For 1K to 4K albedo upsampling, we trained a network (Chen et al., 2023b) on our high-quality dataset, which takes 53 seconds to complete on average. We trained two separate translation networks (Liang et al., 2021) for the specular and displacement maps and the whole translation takes 71 seconds on average. Utilizing these three types of networks, production-quality face assets for an identity can be created in less than 3 minutes on one Nvidia A6000 GPU.
South Asian Female Age 28 Germanic Unisex Age 40
4.2. Results
4.2.1. Qualitative Results
Face Model Generation
Figure 10 shows full assets of 16 identities along with rendered images, created from 8 distinct semantic labels. For each row, two randomly generated results are shown (column 2-5, column 6-9) that follow the same input semantic attributes under different identities. For each subject, the geometry, diffuse albedo map, specular map, and displacement map are created from a set of attributes (1st column) and rendered under three different lighting conditions. These results illustrate the diversity, quality, and effective semantic control of our model.
Texture normalization.
The result of texture normalization is shown as part of Figure 4, where columns 3 and 4 show training data before and after normalization. Skin color is given in column 1. It transforms the UV-space texture with baked-in illumination into albedo textures while preserving the identity of the input texture. Figure 12 shows the effect of skin color control in texture normalization.
Attribute Control.
Figure 13 shows independent control of age and gender in both texture and geometry. In each block, gender varies within each column and age varies within each row, while all other attributes remain fixed. Along the gender axis, we can see that a masculine face generally displays sharper, more pronounced features, including thicker eyebrows. Along the age axis, the most notable difference is the prominence of wrinkles.
Figure 14 shows independent control of skin color in the generator. When skin color is not specified, it is sampled randomly from the skin color distribution of the chosen ethnicity, which is a normal distribution fit to the skin color distribution in the unnormalized training data. The figure presents five skin color samples of each subject, from bright to dark in the range of their ethnicity, while keeping identity and facial features consistent.
Inversion and Editing.
Figures 15 and 16 shows image inversion and editing. Images are first projected into the latent space of the generator by finding the unlabeled code and labeled attributes that generates the closest image using optimization. Edited images are then generated by modifying the labeled attributes while keeping the unlabeled code unchanged.
Additionally, since all generated facial geometries share the same topology from our 3D asset dataset , generic facial geometric attributes editing such as add aging features Figure 13 can be done with some predefined geometry offsets, and blendshapes can be applied to animate the facial assets. For animation examples, please refer to the video.
4.2.2. Comparative Evaluation
Texture normalization.
For a quantitative evaluation, we can take a UV-space image rendered using scanned data, normalize it, and compare the result to the ground truth albedo by calculating the PSNR. We compare our method with FFHQ-UV (Bai et al., 2023). However, in FFHQ-UV, the facial geometry fitting and texture normalization are integrated, and it cannot be applied to UV textures directly. So, we render our scanned data and apply FFHQ-UV. The result is in Table 1.
In addition, on datasets without known ground truth we calculate the Brightness Symmetry Error (BS Error) introduced in FFHQ-UV, as shown in Table 2. The metric measures the dissimilarity of brightness between the left and right halves of the face, based on the assumption that without lighting the brightness of the two halves should be identical. For the Facescape data, we processed it using a method similar to (Li et al., 2020b) to obtain texture data with the same topology. For FFHQ, we employed a 3DMM fitting method to obtain partial textures from the images and then applied the blending described in Section 3.1.2 to complete the textures around the central facial area. For a qualitative evaluation, Figure 9 shows the results of texture normalization compared to FFHQ-UV (Bai et al., 2023). Our normalization results are closer to the target albedo domain, with pure skin color and no lighting baked-in (e.g., highlights). However, the normalized UV generated by FFHQ-UV still retains some of the original illumination, which cannot be used for high-end PBR-base shader as albedo.
| Method | FFHQ-UV | Ours |
| PSNR | 17.64 | 24.67 |
| Method | Facescape* | Light Stage* | FFHQ | FFHQ- UV** | Ours |
| BS Error | 15.2595 | 4.8279 | 28.0723 | 7.293 | 5.8109 |
Facial Attribute Control.
Figure 17 compares our system to DreamFace (Zhang et al., 2023a), a state-of-the-art face asset generation approach. Our system generates real albedo, which reflects true skin color, whereas DreamFace only produces texture under evenly distributed illumination. Also, our generated face model exhibits greater diversity and more closely aligns with the user description as the training dataset of DreamFace mainly consists of Asian people. Compared to general text-to-3D methods like DreamFusion (Poole et al., 2022), LucidDreamer (Liang et al., 2023) and TRELLIS (Xiang et al., 2024), our method produces avatars with impressive detail and production-level textures. DreamFusion creates clear structures, but misses finer skin details. LucidDreamer and TRELLIS output vivid, detailed and realistic faces, but with strong artifacts.
To compare the effectiveness of attribute control between our trained generator and a generic diffusion model (Rombach et al., 2022), we perform age editing on the same subject using both methods, as shown in Figure 19. The portraits generated by the diffusion model, which is the same as the one used in portraits generation, are processed through our data preparation pipeline to obtain final renderings. While diffusion models can achieve age modification through prompt manipulation, our generator produces significantly smoother transitions across age levels. In contrast, diffusion-based results have abrupt feature changes between adjacent slices in the figure. A more detailed comparison of transition smoothness can be found in the supplementary video.
Ours Diffusion
For a quantitative comparison of attribute control of our generator to the other methods, we use CLIP score. We first build a text description of the generated image data for our semantic attributes. We utilize the same ethnicity, gender, and age groups as the attributes we defined for generation to construct descriptors. Following the strategy of (Zhang et al., 2023a), we form text input with the phrase “the realistic face of [DESCRIPTOR]”. After acquiring 5 sets of PBR assets for each descriptor, we render 2500 test images. To calculate the score of Describe3D (Wu et al., 2023), we synthesize facial images by generating the descriptors with the rules mentioned in their work and used the same phrase as input into the CLIP model. Subsequently, we compute the average CLIP score from the ViT-B/16 and ViT-L/14 models. The result, shown in Table 3, highlights the effectiveness of our method in achieving semantic coherence between the generated images and their corresponding text descriptions.
User Study
We conduct a user study to compare the outputs of avatar generation methods, including our proposed method, DreamFace (Zhang et al., 2023a) and Describe3D (Wu et al., 2023). 87 participants rate images synthesized using the corresponding methods in two aspects: description consistency and photorealism. The outcomes in Figure 20 show that our proposed method consistently outperforms the baseline methods in terms of description consistency and photorealism.
4.2.3. Ablation Study
Design of the Texture Normalization Network.
To verify our design for the normalization network, we compare it with the following variations: 1) “uniform patch parameter”: removing the spatial variability of patch parameter by pooling the output of the patch parameter network and broadcasting it to every location, 2) “no color control”: removing external color control and instead using the color calculated from the flat regions of the input image, 3) “Unet”: replacing the normalizer with a UNet, 4) “stride 1 conv net”: replacing the normalizer with a simple convolution network containing only stride-1 convolutions. The results are shown in Figure 21. The network with uniform patch parameters only applies a global color mapping. Due to the invariability of patch parameters it cannot handle strong local variations in lighting, as shown in row 2. Without color control, the quality is comparable, but the model loses its ability to incorporate color information if such information is available. In these examples, the race is known, and the color input is drawn randomly from the color distribution of the race, and the network without color control fails to produce race-accurate output skin color. The UNet enables long-range exchange of information and could modify facial features if that results in a better match between the output normalized texture and the high-quality scan. The scanned data contains a large portion of subjects with heavy eyebrows, and it can be observed that in row 4, the eyebrow appears darker and more expansive in the UNet output than the ground truth. In the stride-1 convnet, as each output pixel depends only on a small patch in the input, there is no coordination across the image, which can result in major failure cases as in row 1. Overall, the ablation study validated that our normalization design performs the best compared to the alternatives.
Effectiveness of Attribute Control.
We evaluate whether, for any given label, the distribution of the features in the images produced by our method closely match the real data distribution. For this purpose, we train a classifier for age, gender and ethnicity, independent from the generator training process. The classifier is trained on 85% of normalized data and tested on the other 15%. We then generate a large number of samples using random codes and attribute labels, and classify these generated samples using the classifier. If the generator controls the attributes accurately, the performance of the classifier should be similar on the test data and on the generated samples. The accuracy (average of all classes) is shown in Table 4, and the confusion matrix is shown in Figure 22. The performance of the classifier on the dataset and the generated samples are similar, and where they differ, the accuracy is generally higher on the generated samples.
| race | gender | age | |
| dataset | |||
| generated | |||
| difference |
| race | gender | age | |
dataset |
generated |
difference |
| generated dataset | |||
Identity Consistency in Attribute Editing.
A key demand for semantic editing of faces is the consistency in facial identity as the attributes of interest (age, gender and ethnicity in our case) are adjusted. Our asset generator is trained to disentangle labeled and unlabeled information and therefore designed to excel in this aspect. However, there is no direct metric proposed in the past that can conveniently evaluate this capability quantitatively: existing face embedding networks for facial identification take all semantic information into consideration, since age, gender and ethnicity are posed as part of one’s identity in these models.
We therefore devise a similarity metric based on linear fitting to evaluate identity preservation of our model in gender and age editing. With the existing face embedding network OpenFace (Amos et al., 2016), we synthesize a large number of random pairs of images, where the two codes for each pair are identical except the controlled attributes. Then, the difference vectors of the embeddings are computed from the two images. The large set of difference vectors are used for linear fitting, where the fitted line should represent the direction corresponding to the controlled attributes (e.g. gender, age) in the embedding space. Note that we exclude ethnicity from this evaluation since its dimensionality is unknown.
The gender-and-age-independent identity similarity can therefore be obtained by projecting the face embeddings onto the complement of the subspace spanned by the estimated gender and age directions. The similarity of two images can be measured by the cosine similarity. We therefore compare the identity similarity of pairs of images where gender and age attributes are edited by our models against random pairs of face images. The results are shown in Table 5, where we find that our model supports attribute editing that maintains a high similarity in identity under our metric. If we change just the age and/or gender, the result should have high similarity to the original, while two random images should have lower similarity.
| change gender | change age | random pair |
Evaluation on the Two-Step Disentanglement Training.
To evaluate the effectiveness of the disentanglement training, we compare our two-stage training design to a typical conditional GAN model as the baseline approach. Examples are shown in Figure 18. We find that our generator can better reproduce unlabeled features (e.g. facial identity): In column 1, the result produced by the baseline GAN has flatter eyebrows; in columns 2 and 4 the baseline result displays less facial hair; in column 3 the baseline result appears more masculine; in column 5 the baseline result shows less wrinkles.
4.2.4. Application
We design an on-the-fly interactive web application based on the proposed method to validate its downstream application value. Figure 23 visualizes the interface for a 3D face asset creation and editing tool. The users are given options to specify ”General Facial Structure” attributes such as age, ethnicity, and gender, and ”Detailed Facial Structure” features including the shape of the face, nose, ears, chin, and more. The web application supports free-viewport manipulation and change of lighting for detailed inspection of the generated results. The application employs a front-end system for rendering and a back-end system for providing services and hosting the web pages. The front-end system uses the Unity game engine (Unity Technologies, 2022) and is built against the WebGL platform. All queries on editing age, ethnicity, and gender are processed by the back-end system, which runs our proposed neural networks to generate the corresponding assets. The mesh object, albedo map, specular map, and displacement map are the major art assets dynamically generated by our models and saved on the cloud storage associated with an ID in our database for later reference. Changes to other facial feature settings are handled locally on the browser, and no new assets are generated by the server. This design eliminates significant delay to ensure an interactive editing experience. In a local session, related assets, such as eye color textures and geometry offset presets, are stored on the back-end server and only loaded on demand. Changes to these modifiable features are updated in the database to allow reproducibility. The back-end system is implemented using ASP.NET Core (Microsoft, 2024) and modularized. Please refer to our supplementary materials for video recordings of the web application.
5. Limitations
While we have made considerable progress towards a production-quality system for semantic-guided generation of PBR face assets, there are aspects of our system that can be improved upon.
By sourcing our training data from diffusion models, we expanded the diversity of our training data greatly. However, our method only works for texture maps, as there is currently no controllable high-quality large-scale generative model for face geometry. While we can refine the geometry with a single-view face reconstruction method, the diversity of the initial geometries is still limited by the scanned dataset.
Our texture normalization matches diffusion-generated UV textures to the scanned-albedo domain. But its accuracy is limited by the small scanned target domain (200 subjects), which can reduce skin-tone fidelity for rare appearances. We therefore estimate a skin color from manually selected flat regions as additional input. With richer scanned-albedo coverage, leveraging general-purpose intrinsic decomposition as additional de-lighting supervision is a potential way to improve normalization further.
Furthermore, the invisible parts of the UV texture projected from the frontal view portraits, generated by the diffusion model, need to be supplemented by the scanned dataset. This requires a well-distributed scan dataset. If the generated portrait has no similar labels in the scanned data, the skin texture may not be accurately completed. Therefore, we aim to explore a multi-view consistent texture synthesis method that can improve our texture completion design.
Since pre-trained diffusion models are used to create the training data, our generation model may inherit any potential biases present in these models. Although we ensure that each attribute has balanced training data through manually defined attributes and distributions in data preparation, it is worth discussing how our method might still inherit issues if the pre-trained models do not express certain attributes diversely. Additionally, due to the high quality of the generated faces, additional care should be taken to prevent misuse and protect privacy.
6. Conclusion
We have introduced a novel 3D face generative model that allows for semantic control and creates high-quality face models, which is unprecedented in 3DMM of human faces. The main contributions of our model are a specifically designed normalization network and a disentangled generator, which can leverage not only high-quality scanned models but also in-the-wild face models, which are abundant in quantity and semantic information. Experiments have demonstrated that our model can effectively normalize faces under arbitrary lighting conditions, generate novel faces, and perform attribute manipulations on the generated 3D face in multiple semantic directions. We believe that the progress our system achieves has great potential for use in many applications, including VFX production, customized digital avatars, and the generation of synthetic training data for other fundamental computer vision research.
Acknowledgements.
This research was sponsored by the U.S. Army Research Laboratory (ARL) under contract number W911NF-14-D-0005. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office, the U.S. Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation.References
- (1)
- 3DScanStore (2023) 3DScanStore. 2023. 3DScanStore. https://www.3dscanstore.com [Online; accessed 24-January-2024].
- Abdal et al. (2019) Rameen Abdal, Yipeng Qin, and Peter Wonka. 2019. Image2StyleGAN: How to Embed Images Into the StyleGAN Latent Space?. In ICCV 2019. 4431–4440.
- Abrevaya et al. (2018) Victoria Fernández Abrevaya, Stefanie Wuhrer, and Edmond Boyer. 2018. Multilinear Autoencoder for 3D Face Model Learning. In 2018 IEEE Winter Conference on Applications of Computer Vision, WACV 2018, Lake Tahoe, NV, USA, March 12-15, 2018. IEEE Computer Society, 1–9. doi:10.1109/WACV.2018.00007
- Amos et al. (2016) Brandon Amos, Bartosz Ludwiczuk, and Mahadev Satyanarayanan. 2016. OpenFace: A general-purpose face recognition library with mobile applications. Technical Report. CMU-CS-16-118, CMU School of Computer Science.
- Bai et al. (2023) Haoran Bai, Di Kang, Haoxian Zhang, Jinshan Pan, and Linchao Bao. 2023. FFHQ-UV: Normalized Facial UV-Texture Dataset for 3D Face Reconstruction. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 362–371.
- Bas et al. (2016) Anil Bas, William A. P. Smith, Timo Bolkart, and Stefanie Wuhrer. 2016. Fitting a 3D Morphable Model to Edges: A Comparison Between Hard and Soft Correspondences. In ACCV 2016 Workshops, Vol. 10117. 377–391.
- Blanz and Vetter (1999) Volker Blanz and Thomas Vetter. 1999. A Morphable Model for the Synthesis of 3D Faces. In Proceedings of the 26th Annual Conference on Computer Graphics and Interactive Techniques, SIGGRAPH 1999, Los Angeles, CA, USA, August 8-13, 1999, Warren N. Waggenspack (Ed.). ACM, 187–194. https://dl.acm.org/citation.cfm?id=311556
- Booth et al. (2018) James Booth, Anastasios Roussos, Allan Ponniah, David Dunaway, and Stefanos Zafeiriou. 2018. Large Scale 3D Morphable Models. Int. J. Comput. Vis. 126, 2-4 (2018), 233–254.
- Booth et al. (2016) James Booth, Anastasios Roussos, Stefanos Zafeiriou, Allan Ponniah, and David Dunaway. 2016. A 3D Morphable Model Learnt from 10, 000 Faces. In 2016 IEEE Conference on Computer Vision and Pattern Recognition, CVPR 2016, Las Vegas, NV, USA, June 27-30, 2016. IEEE Computer Society, 5543–5552. doi:10.1109/CVPR.2016.598
- Burt and Adelson (1983) Peter J Burt and Edward H Adelson. 1983. A multiresolution spline with application to image mosaics. ACM Transactions on Graphics (ToG) 2, 4 (1983), 217–236.
- Cao et al. (2014) Chen Cao, Yanlin Weng, Shun Zhou, Yiying Tong, and Kun Zhou. 2014. FaceWarehouse: A 3D Facial Expression Database for Visual Computing. IEEE Trans. Vis. Comput. Graph. 20, 3 (2014), 413–425. doi:10.1109/TVCG.2013.249
- Cao et al. (2023) Yukang Cao, Yan-Pei Cao, Kai Han, Ying Shan, and Kwan-Yee K Wong. 2023. Dreamavatar: Text-and-shape guided 3d human avatar generation via diffusion models. arXiv preprint arXiv:2304.00916 (2023).
- Chandran et al. (2020) Prashanth Chandran, Derek Bradley, Markus H. Gross, and Thabo Beeler. 2020. Semantic Deep Face Models. In 8th International Conference on 3D Vision, 3DV 2020, Virtual Event, Japan, November 25-28, 2020, Vitomir Struc and Francisco Gómez Fernández (Eds.). IEEE, 345–354.
- Chen et al. (2019) Anpei Chen, Zhang Chen, Guli Zhang, Kenny Mitchell, and Jingyi Yu. 2019. Photo-realistic facial details synthesis from single image. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 9429–9439.
- Chen et al. (2023a) Rui Chen, Yongwei Chen, Ningxin Jiao, and Kui Jia. 2023a. Fantasia3d: Disentangling geometry and appearance for high-quality text-to-3d content creation. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 22246–22256.
- Chen et al. (2024) Xi Chen, Sida Peng, Dongchen Yang, Yuan Liu, Bowen Pan, Chengfei Lv, and Xiaowei Zhou. 2024. Intrinsicanything: Learning diffusion priors for inverse rendering under unknown illumination. In European Conference on Computer Vision. Springer, 450–467.
- Chen et al. (2023b) Zheng Chen, Yulun Zhang, Jinjin Gu, Linghe Kong, Xiaokang Yang, and Fisher Yu. 2023b. Dual Aggregation Transformer for Image Super-Resolution. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 12312–12321.
- Dai et al. (2020) Hang Dai, Nick E. Pears, William A. P. Smith, and Christian Duncan. 2020. Statistical Modeling of Craniofacial Shape and Texture. Int. J. Comput. Vis. 128, 2 (2020), 547–571. doi:10.1007/s11263-019-01260-7
- Debevec et al. (2000) Paul Debevec, Tim Hawkins, Chris Tchou, Haarm-Pieter Duiker, Westley Sarokin, and Mark Sagar. 2000. Acquiring the reflectance field of a human face. In Proceedings of the 27th annual conference on Computer graphics and interactive techniques. 145–156.
- Dib et al. (2021) Abdallah Dib, Cedric Thebault, Junghyun Ahn, Philippe-Henri Gosselin, Christian Theobalt, and Louis Chevallier. 2021. Towards high fidelity monocular face reconstruction with rich reflectance using self-supervised learning and ray tracing. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 12819–12829.
- Egger et al. (2020) Bernhard Egger, William A. P. Smith, Ayush Tewari, Stefanie Wuhrer, Michael Zollhöfer, Thabo Beeler, Florian Bernard, Timo Bolkart, Adam Kortylewski, Sami Romdhani, Christian Theobalt, Volker Blanz, and Thomas Vetter. 2020. 3D Morphable Face Models - Past, Present, and Future. ACM Trans. Graph. 39, 5 (2020), 157:1–157:38.
- Fadaeinejad et al. (2025) Amin Fadaeinejad, Abdallah Dib, Luiz Gustavo Hafemann, Emeline Got, Trevor Anderson, Amaury Depierre, Nikolaus F Troje, Marcus A Brubaker, and Marc-André Carbonneau. 2025. Geometry-Aware Texture Generation for 3D Head Modeling with Artist-driven Control. In Proceedings of the Computer Vision and Pattern Recognition Conference. 6196–6206.
- Fang et al. (2021) Zhixin Fang, Libai Cai, and Gang Wang. 2021. MetaHuman Creator The starting point of the metaverse. In 2021 International Symposium on Computer Technology and Information Science (ISCTIS). IEEE, 154–157.
- Gecer et al. (2019) Baris Gecer, Stylianos Ploumpis, Irene Kotsia, and Stefanos Zafeiriou. 2019. Ganfit: Generative adversarial network fitting for high fidelity 3d face reconstruction. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 1155–1164.
- Ghafourzadeh et al. (2020) Donya Ghafourzadeh, Cyrus Rahgoshay, Sahel Fallahdoust, Andre Beauchamp, Adeline Aubame, Tiberiu Popa, and Eric Paquette. 2020. Part-Based 3D Face Morphable Model with Anthropometric Local Control. In Proceedings of the 45th Graphics Interface Conference 2020, Toronto, ON, Canada, May 28-29, 2020, David I. W. Levin, Fanny Chevalier, and Alec Jacobson (Eds.). Canadian Human-Computer Communications Society, 7–16.
- Ghosh et al. (2011) Abhijeet Ghosh, Graham Fyffe, Borom Tunwattanapong, Jay Busch, Xueming Yu, and Paul Debevec. 2011. Multiview face capture using polarized spherical gradient illumination. ACM Transactions on Graphics (TOG) 30, 6 (2011), 1–10.
- Goodfellow et al. (2014) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2014. Generative Adversarial Nets. In NeurIPS 2014. 2672–2680.
- Härkönen et al. (2020) Erik Härkönen, Aaron Hertzmann, Jaakko Lehtinen, and Sylvain Paris. 2020. GANSpace: Discovering Interpretable GAN Controls. In Advances in Neural Information Processing Systems 33: Annual Conference on Neural Information Processing Systems 2020, NeurIPS 2020, December 6-12, 2020, virtual, Hugo Larochelle, Marc’Aurelio Ranzato, Raia Hadsell, Maria-Florina Balcan, and Hsuan-Tien Lin (Eds.). https://proceedings.neurips.cc/paper/2020/hash/6fe43269967adbb64ec6149852b5cc3e-Abstract.html
- Ho and Salimans (2022) Jonathan Ho and Tim Salimans. 2022. Classifier-free diffusion guidance. arXiv preprint arXiv:2207.12598 (2022).
- Hong et al. (2022) Fangzhou Hong, Mingyuan Zhang, Liang Pan, Zhongang Cai, Lei Yang, and Ziwei Liu. 2022. Avatarclip: Zero-shot text-driven generation and animation of 3d avatars. arXiv preprint arXiv:2205.08535 (2022).
- Huynh et al. (2018) Loc Huynh, Weikai Chen, Shunsuke Saito, Jun Xing, Koki Nagano, Andrew Jones, Paul Debevec, and Hao Li. 2018. Mesoscopic facial geometry inference using deep neural networks. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition. 8407–8416.
- Jiang et al. (2021) Yuming Jiang, Ziqi Huang, Xingang Pan, Chen Change Loy, and Ziwei Liu. 2021. Talk-to-Edit: Fine-Grained Facial Editing via Dialog. In 2021 IEEE/CVF International Conference on Computer Vision, ICCV 2021, Montreal, QC, Canada, October 10-17, 2021. IEEE, 13779–13788.
- Karras et al. (2020) Tero Karras, Samuli Laine, Miika Aittala, Janne Hellsten, Jaakko Lehtinen, and Timo Aila. 2020. Analyzing and improving the image quality of stylegan. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8110–8119.
- Khodadadeh et al. (2022) Siavash Khodadadeh, Shabnam Ghadar, Saeid Motiian, Wei-An Lin, Ladislau Bölöni, and Ratheesh Kalarot. 2022. Latent to Latent: A Learned Mapper for Identity Preserving Editing of Multiple Face Attributes in StyleGAN-generated Images. In IEEE/CVF Winter Conference on Applications of Computer Vision, WACV 2022, Waikoloa, HI, USA, January 3-8, 2022. IEEE, 3677–3685.
- Kocsis et al. (2024) Peter Kocsis, Vincent Sitzmann, and Matthias Nießner. 2024. Intrinsic image diffusion for indoor single-view material estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 5198–5208.
- Kowalski et al. (2020) Marek Kowalski, Stephan J. Garbin, Virginia Estellers, Tadas Baltrusaitis, Matthew Johnson, and Jamie Shotton. 2020. CONFIG: Controllable Neural Face Image Generation. In Computer Vision - ECCV 2020 - 16th European Conference, Glasgow, UK, August 23-28, 2020, Proceedings, Part XI (Lecture Notes in Computer Science, Vol. 12356), Andrea Vedaldi, Horst Bischof, Thomas Brox, and Jan-Michael Frahm (Eds.). Springer, 299–315.
- Laine (2018) Samuli Laine. 2018. Feature-Based Metrics for Exploring the Latent Space of Generative Models. In 6th International Conference on Learning Representations, ICLR 2018, Vancouver, BC, Canada, April 30 - May 3, 2018, Workshop Track Proceedings. OpenReview.net. https://openreview.net/forum?id=BJslDBkwG
- Lattas et al. (2022) Alexandros Lattas, Yiming Lin, Jayanth Kannan, Ekin Ozturk, Luca Filipi, Giuseppe Claudio Guarnera, Gaurav Chawla, and Abhijeet Ghosh. 2022. Practical and scalable desktop-based high-quality facial capture. In European Conference on Computer Vision. Springer, 522–537.
- Lattas et al. (2020) Alexandros Lattas, Stylianos Moschoglou, Baris Gecer, Stylianos Ploumpis, Vasileios Triantafyllou, Abhijeet Ghosh, and Stefanos Zafeiriou. 2020. AvatarMe: Realistically Renderable 3D Facial Reconstruction” in-the-wild”. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 760–769.
- Lattas et al. (2023) Alexandros Lattas, Stylianos Moschoglou, Stylianos Ploumpis, Baris Gecer, Jiankang Deng, and Stefanos Zafeiriou. 2023. FitMe: Deep Photorealistic 3D Morphable Model Avatars. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8629–8640.
- Lattas et al. (2021) Alexandros Lattas, Stylianos Moschoglou, Stylianos Ploumpis, Baris Gecer, Abhijeet Ghosh, and Stefanos Zafeiriou. 2021. Avatarme++: Facial shape and brdf inference with photorealistic rendering-aware gans. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 12 (2021), 9269–9284.
- Li et al. (2020c) Jiaman Li, Zhengfei Kuang, Yajie Zhao, Mingming He, Karl Bladin, and Hao Li. 2020c. Dynamic facial asset and rig generation from a single scan. ACM Transactions on Graphics (TOG) 39, 6 (2020), 1–18.
- Li et al. (2020d) Jiaman Li, Zhengfei Kuang, Yajie Zhao, Mingming He, Karl Bladin, and Hao Li. 2020d. Dynamic facial asset and rig generation from a single scan. ACM Trans. Graph. 39, 6 (2020), 215:1–215:18.
- Li et al. (2020b) Ruilong Li, Karl Bladin, Yajie Zhao, Chinmay Chinara, Owen Ingraham, Pengda Xiang, Xinglei Ren, Pratusha Prasad, Bipin Kishore, Jun Xing, et al. 2020b. Learning formation of physically-based face attributes. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 3410–3419.
- Li et al. (2020a) Ruilong Li, Karl Bladin, Yajie Zhao, Chinmay Chinara, Owen Ingraham, Pengda Xiang, Xinglei Ren, Pratusha Prasad, Bipin Kishore, Jun Xing, and Hao Li. 2020a. Learning Formation of Physically-Based Face Attributes. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 3407–3416. doi:10.1109/CVPR42600.2020.00347
- Li et al. (2017) Tianye Li, Timo Bolkart, Michael J. Black, Hao Li, and Javier Romero. 2017. Learning a model of facial shape and expression from 4D scans. ACM Trans. Graph. 36, 6 (2017), 194:1–194:17. doi:10.1145/3130800.3130813
- Li et al. (2025) Zhibing Li, Tong Wu, Jing Tan, Mengchen Zhang, Jiaqi Wang, and Dahua Lin. 2025. IDArb: Intrinsic Decomposition for Arbitrary Number of Input Views and Illuminations. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=uuef1HP6X7
- Liang et al. (2021) Jie Liang, Hui Zeng, and Lei Zhang. 2021. High-resolution photorealistic image translation in real-time: A laplacian pyramid translation network. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 9392–9400.
- Liang et al. (2023) Yixun Liang, Xin Yang, Jiantao Lin, Haodong Li, Xiaogang Xu, and Yingcong Chen. 2023. LucidDreamer: Towards High-Fidelity Text-to-3D Generation via Interval Score Matching. arXiv:2311.11284 [cs.CV]
- Lin et al. (2023) Chen-Hsuan Lin, Jun Gao, Luming Tang, Towaki Takikawa, Xiaohui Zeng, Xun Huang, Karsten Kreis, Sanja Fidler, Ming-Yu Liu, and Tsung-Yi Lin. 2023. Magic3d: High-resolution text-to-3d content creation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 300–309.
- Liu et al. (2022) Shichen Liu, Yunxuan Cai, Haiwei Chen, Yichao Zhou, and Yajie Zhao. 2022. Rapid Face Asset Acquisition with Recurrent Feature Alignment. ACM Trans. Graph. 41, 6, Article 214 (nov 2022), 17 pages. doi:10.1145/3550454.3555509
- Luo et al. (2024) Jundan Luo, Duygu Ceylan, Jae Shin Yoon, Nanxuan Zhao, Julien Philip, Anna Frühstück, Wenbin Li, Christian Richardt, and Tuanfeng Wang. 2024. Intrinsicdiffusion: Joint intrinsic layers from latent diffusion models. In ACM SIGGRAPH 2024 Conference Papers. 1–11.
- Metzer et al. (2023) Gal Metzer, Elad Richardson, Or Patashnik, Raja Giryes, and Daniel Cohen-Or. 2023. Latent-nerf for shape-guided generation of 3d shapes and textures. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 12663–12673.
- Michel et al. (2022) Oscar Michel, Roi Bar-On, Richard Liu, Sagie Benaim, and Rana Hanocka. 2022. Text2mesh: Text-driven neural stylization for meshes. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 13492–13502.
- Microsoft (2024) Microsoft. 2024. ASP.NET Core. https://dotnet.microsoft.com/apps/aspnet Accessed: 2024-05-19.
- Nagano et al. (2018) Koki Nagano, Jaewoo Seo, Jun Xing, Lingyu Wei, Zimo Li, Shunsuke Saito, Aviral Agarwal, Jens Fursund, and Hao Li. 2018. paGAN: real-time avatars using dynamic textures. ACM Transactions on Graphics (TOG) 37, 6 (2018), 1–12.
- Paysan et al. (2009) Pascal Paysan, Reinhard Knothe, Brian Amberg, Sami Romdhani, and Thomas Vetter. 2009. A 3D Face Model for Pose and Illumination Invariant Face Recognition. In Sixth IEEE International Conference on Advanced Video and Signal Based Surveillance, AVSS 2009, 2-4 September 2009, Genova, Italy. 296–301.
- Poole et al. (2022) Ben Poole, Ajay Jain, Jonathan T Barron, and Ben Mildenhall. 2022. DreamFusion: Text-to-3D using 2D Diffusion. In The Eleventh International Conference on Learning Representations.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, Gretchen Krueger, and Ilya Sutskever. 2021. Learning Transferable Visual Models From Natural Language Supervision. arXiv:2103.00020 [cs.CV]
- Radford et al. (2016) Alec Radford, Luke Metz, and Soumith Chintala. 2016. Unsupervised Representation Learning with Deep Convolutional Generative Adversarial Networks. In ICLR 2016.
- Ranjan et al. (2018) Anurag Ranjan, Timo Bolkart, Soubhik Sanyal, and Michael J. Black. 2018. Generating 3D Faces Using Convolutional Mesh Autoencoders. In Computer Vision - ECCV 2018 - 15th European Conference, Munich, Germany, September 8-14, 2018, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 11207), Vittorio Ferrari, Martial Hebert, Cristian Sminchisescu, and Yair Weiss (Eds.). Springer, 725–741. doi:10.1007/978-3-030-01219-9_43
- Riviere et al. (2020) Jérémy Riviere, Paulo Gotardo, Derek Bradley, Abhijeet Ghosh, and Thabo Beeler. 2020. Single-shot high-quality facial geometry and skin appearance capture. (2020).
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 10684–10695.
- Saito et al. (2017) Shunsuke Saito, Lingyu Wei, Liwen Hu, Koki Nagano, and Hao Li. 2017. Photorealistic facial texture inference using deep neural networks. In Proceedings of the IEEE conference on computer vision and pattern recognition. 5144–5153.
- Shao et al. (2018) Hang Shao, Abhishek Kumar, and P. Thomas Fletcher. 2018. The Riemannian Geometry of Deep Generative Models. In 2018 IEEE Conference on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2018, Salt Lake City, UT, USA, June 18-22, 2018. IEEE Computer Society, 315–323. doi:10.1109/CVPRW.2018.00071
- Shen et al. (2020) Yujun Shen, Jinjin Gu, Xiaoou Tang, and Bolei Zhou. 2020. Interpreting the Latent Space of GANs for Semantic Face Editing. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 9240–9249. doi:10.1109/CVPR42600.2020.00926
- Shen and Zhou (2020) Yujun Shen and Bolei Zhou. 2020. Closed-Form Factorization of Latent Semantics in GANs. CoRR abs/2007.06600 (2020). arXiv:2007.06600 https://arxiv.org/abs/2007.06600
- Smith et al. (2020) William A. P. Smith, Alassane Seck, Hannah Dee, Bernard Tiddeman, Joshua B. Tenenbaum, and Bernhard Egger. 2020. A Morphable Face Albedo Model. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 5010–5019. doi:10.1109/CVPR42600.2020.00506
- Tewari et al. (2020a) Ayush Tewari, Mohamed Elgharib, Gaurav Bharaj, Florian Bernard, Hans-Peter Seidel, Patrick Pérez, Michael Zollhöfer, and Christian Theobalt. 2020a. StyleRig: Rigging StyleGAN for 3D Control Over Portrait Images. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. Computer Vision Foundation / IEEE, 6141–6150.
- Tewari et al. (2020b) Ayush Tewari, Mohamed Elgharib, Mallikarjun B. R., Florian Bernard, Hans-Peter Seidel, Patrick Pérez, Michael Zollhöfer, and Christian Theobalt. 2020b. PIE: portrait image embedding for semantic control. ACM Trans. Graph. 39, 6 (2020), 223:1–223:14.
- Thies et al. (2016) Justus Thies, Michael Zollhöfer, Marc Stamminger, Christian Theobalt, and Matthias Nießner. 2016. Face2Face: Real-Time Face Capture and Reenactment of RGB Videos. In CVPR 2016. 2387–2395.
- Triplegangers (2021) Triplegangers. 2021. Triplegangers Face Models. https://triplegangers.com/. Online; Accessed: 2021-12-05.
- Unity Technologies (2022) Unity Technologies. 2022. Unity Game Engine. https://unity.com/ Accessed: 2024-05-19.
- Upchurch et al. (2017) Paul Upchurch, Jacob R. Gardner, Geoff Pleiss, Robert Pless, Noah Snavely, Kavita Bala, and Kilian Q. Weinberger. 2017. Deep Feature Interpolation for Image Content Changes. In 2017 IEEE Conference on Computer Vision and Pattern Recognition, CVPR 2017, Honolulu, HI, USA, July 21-26, 2017. IEEE Computer Society, 6090–6099.
- Wang et al. (2022) Can Wang, Menglei Chai, Mingming He, Dongdong Chen, and Jing Liao. 2022. Cross-Domain and Disentangled Face Manipulation with 3D Guidance. IEEE Transactions on Visualization and Computer Graphics (2022).
- Wu et al. (2023) Menghua Wu, Hao Zhu, Linjia Huang, Yiyu Zhuang, Yuanxun Lu, and Xun Cao. 2023. High-fidelity 3D face generation from natural language descriptions. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 4521–4530.
- Xiang et al. (2024) Jianfeng Xiang, Zelong Lv, Sicheng Xu, Yu Deng, Ruicheng Wang, Bowen Zhang, Dong Chen, Xin Tong, and Jiaolong Yang. 2024. Structured 3D Latents for Scalable and Versatile 3D Generation. arXiv preprint arXiv:2412.01506 (2024).
- Xiang et al. (2021) Sitao Xiang, Yuming Gu, Pengda Xiang, Menglei Chai, Hao Li, Yajie Zhao, and Mingming He. 2021. DisUnknown: Distilling Unknown Factors for Disentanglement Learning. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 14810–14819.
- Yamaguchi et al. (2018) Shugo Yamaguchi, Shunsuke Saito, Koki Nagano, Yajie Zhao, Weikai Chen, Kyle Olszewski, Shigeo Morishima, and Hao Li. 2018. High-fidelity facial reflectance and geometry inference from an unconstrained image. ACM Transactions on Graphics (TOG) 37, 4 (2018), 1–14.
- Yang et al. (2020) Haotian Yang, Hao Zhu, Yanru Wang, Mingkai Huang, Qiu Shen, Ruigang Yang, and Xun Cao. 2020. FaceScape: a large-scale high quality 3D face Dataset and detailed riggable 3D face prediction. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 601–610.
- Zhang (2023) Lvmin Zhang. 2023. ControlNet v1.1. https://huggingface.co/lllyasviel/ControlNet-v1-1/. Accessed: 2024-05-19.
- Zhang et al. (2023a) Longwen Zhang, Qiwei Qiu, Hongyang Lin, Qixuan Zhang, Cheng Shi, Wei Yang, Ye Shi, Sibei Yang, Lan Xu, and Jingyi Yu. 2023a. DreamFace: Progressive Generation of Animatable 3D Faces under Text Guidance. arXiv preprint arXiv:2304.03117 (2023).
- Zhang et al. (2023b) Lvmin Zhang, Anyi Rao, and Maneesh Agrawala. 2023b. Adding Conditional Control to Text-to-Image Diffusion Models.
- Zheng et al. (2021) Yutong Zheng, Yu-Kai Huang, Ran Tao, Zhiqiang Shen, and Marios Savvides. 2021. Unsupervised Disentanglement of Linear-Encoded Facial Semantics. In IEEE Conference on Computer Vision and Pattern Recognition, CVPR 2021, virtual, June 19-25, 2021. Computer Vision Foundation / IEEE, 3917–3926.
- Zhou et al. (2024a) Mingyuan Zhou, Rakib Hyder, Ziwei Xuan, and Guojun Qi. 2024a. UltrAvatar: A Realistic Animatable 3D Avatar Diffusion Model with Authenticity Guided Textures. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 1238–1248.
- Zhou et al. (2024b) Zhenglin Zhou, Fan Ma, Hehe Fan, Zongxin Yang, and Yi Yang. 2024b. HeadStudio: Text to Animatable Head Avatars with 3D Gaussian Splatting. In ECCV.
- Zhu et al. (2020) Jiapeng Zhu, Yujun Shen, Deli Zhao, and Bolei Zhou. 2020. In-Domain GAN Inversion for Real Image Editing. In ECCV 2020, Vol. 12362. 592–608.
Appendix A Details of the Comparisons and User Study
Six short textual prompts, each describing a distinct subject identity in terms of age, gender, and ethnicity, were used to generate 3D human models with each method. These prompts served as the primary input to the generation process and were also shown to participants during evaluation. The full list of prompts included: “Elderly Asian Female,” “Young Adult Western Female,” “Teenage African Male,” “Young Western Male,” “Young Asian Female,” and “Middle-Aged African Male.” All other inputs and settings followed the default configurations or recommended usage described in the original implementations of DreamFace (Zhang et al., 2023a) and Describe3D (Wu et al., 2023).
For each prompt, participants viewed the outputs generated by three different methods. The three outputs were displayed in randomized order. The sequence of prompts was also randomized for each participant.
Each output was rated using two questions on a five-point Likert scale (Very Poor to Excellent): (1) “How well does the model match the description?” and (2) “How would you rate the skin detail quality?” In a small number of trials, a third question was included as an attention check, instructing participants to select a specific response (e.g., “Please select ’Poor’.”). Participants who failed one or more of these checks were excluded from the analysis.
Appendix B Hair and Accessories
To support more complete and realistic digital head models, our system includes hair and common accessories such as eyeglasses, hats, and masks. Hair is modeled using Maya XGen and is attached to a predefined scalp mesh. Since all head models share the same topology, the hair can be transferred across different identities by deforming it according to the shared vertex correspondence.
As for accessories, eyeglasses are rigged individually, and hats and masks are deformed using lattice-based transformations, allowing them to fit varying face geometries.
Appendix C Challenging Cases
The examples in Figure 24 demonstrate that directly processing images generated by general-purpose diffusion models to UV space often leads to significant artifacts, such as malformed teeth, unnatural tone, and local geometric distortions. Some are due to imperfect control over the diffusion model, where the final synthesized image does not align pixel-perfectly with the input geometry, causing mismatches in the UV projection. Others are inherent to the generative process. The model produces undesired color tones due to a lack of structural constraints. Although we manually filtered out abnormal textures when constructing 44,000 UV textures, these issues highlight the need for a generative model specifically trained for texture synthesis.
Appendix D Dataset demographic distribution
The Figure 26 illustrates the distribution of demographic information in our high-quality scanned dataset.